OpenStudy (linyu):

find volume of the solid generated by revolving the area between the curve y=cosx/x and the x-axis for pi/6≤x≤pi/2 about the y-axis.

OpenStudy (anonymous):

this must be solved by "Sheel method" to find the volumes

OpenStudy (linyu):

can you give me the source to learn?

OpenStudy (anonymous):

\[ V=2\pi\int_a^b r(x)y(x)dx \] where r(x) = radius of the shell.. how far is the shell from the axis of rotation y(x) = height of the shell dx = thickness of the shell

OpenStudy (anonymous):

now, here we find the shell at a distance of "x" from the axis of rotation the height of the shell at that location is: \(y={\cos x\over x}\) the limits are from \({\pi\over6}\le x\le{\pi\over2}\)

OpenStudy (linyu):

I have no clue to solve this. Please give me a way to learn.

OpenStudy (anonymous):